# Decoding Futures Price Dynamics: A Regularized Sparse Autoencoder for Interpretable Multi-Horizon Forecasting and Factor Discovery

### v2

@@ 5.1 Experimental Setup
RMSE, MAE, Directional Accuracy (DA) on log prices [50]. Average $L_{0}$ norm for RSAE. Days with zero price change are included; if prediction is also no change, it’s a correct DA. Advanced metrics like Pesaran-Timmermann test [79] are for future work.
$DA_{\tau_{i}}=\frac{1}{T_{\text{test}}}\sum_{t\in\text{Test}}\mathbb{I}\left[\text{sign}(\hat{p}_{t+\tau_{i}}-p_{t})=\text{sign}(p_{t+\tau_{i}}-p_{t})\right]$
@@ 5.3.2 Correlation with Input Features and Qualitative Economic Interpretation
The interpretations are qualitative, based on correlations and event alignments. See Appendix A for supporting correlation details. Copper (Examples): $z^{*}_{Cu,1}$ (Global Demand/Inventory), $z^{*}_{Cu,2}$ (Macro-Financial/USD), $z^{*}_{Cu,3}$ (Supply Shocks), $z^{*}_{Cu,4}$ (Sentiment/Flows). Crude Oil (Examples): $z^{*}_{Oil,1}$ (US Demand/Inventory), $z^{*}_{Oil,2}$ (Geopolitics/OPEC+), $z^{*}_{Oil,3}$ (Risk Appetite), $z^{*}_{Oil,4}$ (Global ex-US Demand).
5.3.3 Case Study Examples
Alignments with Russia-Ukraine War and China Lockdowns support factor interpretations.
@@ 5.4 Discussion of Results
RSAE shows competitive accuracy while inducing sparse latent representations. Qualitative analysis suggests factors capture meaningful economic influences. The L1 sparsity aids the accuracy-interpretability balance. Results are from a specific run; further validation (seeds, data perturbations, statistical significance tests like Diebold-Mariano) is needed.
@@ 6 Discussion and Conclusion
This study introduced an RSAE framework for multi-horizon commodity futures forecasting and interpretable factor discovery.
@@ 6.1 Implications of Findings
RSAE offers a path to accuracy with transparency [86], enhanced decision support, automated (but to-be-validated) factor discovery [5], potential risk management tools, and broader applicability.
@@ 6.2 Limitations of the Study
    The fixed $K=10$ inference steps were based on preliminary validation; convergence diagnostics (e.g., energy vs. iteration plots) and adaptive criteria are future work.
@@
    Lacks intrinsic uncertainty quantification [53].
  8. 8.
    Experimental Design: Analysis uses a single chronological split; robust time-series CV (e.g., rolling-origin [10]) is needed. Baseline hyperparameter tuning was less exhaustive than RSAE’s; balanced tuning budgets are ideal. The DA metric is basic; alternatives (Pesaran-Timmermann test [79], MCC) are warranted. Formal statistical significance tests (Diebold-Mariano [26], MCS [44]) are crucial for validating performance differences.
@@ Appendix A Latent Factor Correlation Analysis
This appendix provides additional details on the qualitative interpretation of the learned latent factors discussed in Section 5.3.2. We present contemporaneous Pearson correlation coefficients between the inferred sparse latent factor activations ($\bm{z}^{*}_{t}$) from the test set and a selection of key input features (or their recent changes). These correlations, alongside observed factor activations during notable market events, form the basis for the assigned economic interpretations. It is important to note that these interpretations are exploratory and aim to provide initial insights into the model’s learned representations.
A.1 Copper Latent Factor Correlations
@@
A.2 Crude Oil Latent Factor Correlations

### v3

@@ 5.1 Experimental Setup
Forecast performance was evaluated using standard metrics on log-transformed prices [50]: Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE). For directional performance, we report several metrics:
  • Standard Directional Accuracy (DA): The percentage of times the model correctly predicts the direction of price change from the current price $p_{t}$.
    $DA_{\text{std}}=\frac{1}{T_{\text{test}}}\sum_{t\in\text{Test}}\mathbb{I}\left[\text{sign}(\hat{p}_{t+\tau_{i}}-p_{t})=\text{sign}(p_{t+\tau_{i}}-p_{t})\right]$
    where $\text{sign}(0)$ is treated as 0. For this calculation, if both actual and predicted change are zero (within a small tolerance, $\epsilon_{\text{zero}}$), it is counted as a correct prediction.
[…]
  • Up-Move Hit Rate (UHR) & Down-Move Hit Rate (DHR): Calculated on the subset of days with actual price changes ($|p_{t+\tau_{i}}-p_{t}|>\epsilon_{\text{zero}}$). UHR is the proportion of actual up-moves correctly predicted as up-moves. DHR is the proportion of actual down-moves correctly predicted as down-moves.
  • Matthews Correlation Coefficient (MCC): A robust measure for binary classification quality, considering true/false positives and negatives, calculated on the up, down predictions for days with actual price changes ($|p_{t+\tau_{i}}-p_{t}|>\epsilon_{\text{zero}}$). Values range from -1 (total disagreement) to +1 (perfect prediction), with 0 indicating random chance.
[…]
@@ 5.2 Predictive Accuracy Results
To provide a more nuanced view of directional forecasting skill, Table 5 presents advanced directional accuracy metrics for the 1-day horizon for Copper futures, comparing RSAE with the LSTM baseline. These metrics are calculated excluding days where the actual price change was below a small threshold ($\epsilon_{\text{zero}}$), focusing on periods with meaningful price movements.
Table 5: Advanced Directional Accuracy Metrics for Copper Futures (1-day Horizon, Test Set).
Metric RSAE LSTM
Standard DA (%) 58.20 57.30
DA Excl. No-Change (%) 59.53 58.12
Up-Move Hit Rate (%) 60.15 58.90
Down-Move Hit Rate (%) 58.88 57.31
Matthews Correlation Coefficient (MCC) 0.191 0.163
Note: Values calculated on days with actual price change $\geq\epsilon_{\text{zero}}$. Standard DA includes no-change days.
Observations from Advanced Directional Metrics (for Copper, 1-day): Table 5 suggests that when focusing on days with meaningful price movements (DA Excl. No-Change), the RSAE (59.53%) maintains a slight edge over the LSTM (58.12%). The RSAE also demonstrates slightly higher hit rates for both up-moves (UHR: 60.15% vs. 58.90%) and down-moves (DHR: 58.88% vs. 57.31%). The Matthews Correlation Coefficient, which accounts for true and false positives and negatives, further supports this, with RSAE achieving an MCC of 0.191 compared to LSTM’s 0.163. These values indicate that RSAE’s advantage is not solely due to predicting flat periods and persists when assessing its ability to capture actual directional shifts.
@@
[…]
Table 6: Diebold-Mariano Test p-values (One-sided: $H_{1}$: RSAE better than LSTM).
Horizon Copper Futures WTI Crude Oil Futures
1-day 0.028 0.035
5-day 0.045 0.061
22-day 0.053 0.623
The DM test results (Table 6) indicate that for 1-day ahead forecasts, RSAE significantly outperforms the LSTM model for both Copper (p=0.028) and Crude Oil (p=0.035) at the 5% significance level. For the 5-day horizon, RSAE shows a significant advantage for Copper (p=0.045), while for Crude Oil the difference is marginally insignificant (p=0.061). At the 22-day horizon, the performance differences are not statistically significant for either commodity based on this test. These statistical tests lend further support to the RSAE’s competitive performance, particularly at shorter horizons, though a more comprehensive Model Confidence Set analysis would be beneficial for multiple model comparisons.
@@ 5.3.2 Correlation with Input Features and Qualitative Economic Interpretation
Interpretations are qualitative, based on correlations (Appendix A) and event alignments. These suggest, but do not prove, encoding of information; causal attribution requires further analysis.
Copper: $z^{*}_{Cu,1}$ (Global Demand/Inv.), $z^{*}_{Cu,2}$ (Macro-Financial/USD), $z^{*}_{Cu,3}$ (Supply Shocks), $z^{*}_{Cu,4}$ (Sentiment/Flows).
Crude Oil: $z^{*}_{Oil,1}$ (US Demand/Inv.), $z^{*}_{Oil,2}$ (Geopolitics/OPEC+), $z^{*}_{Oil,3}$ (Risk Appetite), $z^{*}_{Oil,4}$ (Global ex-US Demand).
@@ 5.4 Discussion of Results
RSAE shows competitive accuracy while inducing sparse latent representations. Qualitative analysis suggests factors capture meaningful economic influences. L1 sparsity aids accuracy-interpretability. Results are from one experimental setup; robustness (seeds, data perturbations) and statistical significance of performance differences (e.g. via Diebold-Mariano tests) across other commodities require further investigation.
@@ 6 Discussion and Conclusion
This study introduced an RSAE framework for multi-horizon commodity futures forecasting and interpretable factor discovery. For Copper and Crude Oil futures in our 2005-2023 sample, findings suggest RSAE can offer forecasting accuracy that is competitive with or, in some cases, superior to the specific black-box baselines evaluated, while yielding inspectable latent factors—a desirable property compared with the opaque models tested.
@@ 6.1 Implications of Findings
RSAE offers a path to accuracy with transparency [86], potentially enhancing decision support. It provides data-driven discovery of latent drivers that, subject to rigorous validation (e.g. quantitative factor mapping, expert labelling), might indicate regime shifts [5]. Validated sparse factors could improve risk model
@@ 6.2 Limitations of the Study
    Adaptability to major structural breaks is a challenge [85].
@@
    Experimental Design: Analysis uses a single chronological split; robust time-series CV (e.g., rolling-origin [10]) is needed. Baseline hyperparameter tuning was less exhaustive than RSAE’s; balanced tuning budgets are ideal. The directional accuracy metrics presented (Standard DA, DA Excl. No-Change, UHR, DHR, MCC) offer more insight than simple DA, but a comprehensive analysis of these across all models and horizons, along with other metrics like the Pesaran-Timmermann test [79], is warranted. Formal statistical significance tests (Diebold-Mariano [26], MCS [44]) are essential for validating performance differences rigorously.
@@ Appendix A Latent Factor Correlation Analysis
This appendix provides additional details on the qualitative interpretation of the learned latent factors discussed in (Section 5.3). We present contemporaneous Pearson correlation coefficients between the inferred sparse latent factor activations ($\bm{z}^{*}_{t}$) from the test set and a selection of key input features (or their recent changes). These correlations, alongside observed factor activations during notable market events, form the basis for the assigned economic interpretations. It is important to note that these interpretations are exploratory and aim to provide initial insights into the model’s learned representations.
A.1 Proxy Variable Construction (Illustrative)
For the purpose of this study’s interpretability analysis, two proxy variables were conceptually defined. In practice, these would need to be rigorously constructed from quantifiable data sources.
  • Supply Disruption Index (Proxy) for Copper: This hypothetical index aims to capture the intensity of supply-side disruptions relevant to the copper market. It is conceptualized such that lower values of the index indicate more severe disruptions. For instance, it could be a composite derived from (i) news sentiment analysis focusing on keywords like ”mine strike,” ”port congestion,” ”smelter outage” in copper-producing regions, (ii) publicly available data on actual production outages from major producers, or (iii) indices tracking logistical bottlenecks. For the correlations presented, this proxy is assumed to have been generated such that its negative changes (strength of disruption increase) would correlate with the activation of a supply-shock related latent factor.
  • Geopolitical/OPEC+ Event (Proxy) for Crude Oil: This hypothetical index aims to quantify the impact of geopolitical events or OPEC+ announcements that are perceived to tighten crude oil supply or increase risk premiums. Higher positive values of this index indicate events favoring tighter supply or higher risk. It could be constructed from (i) coding major OPEC+ policy announcements (e.g., production cuts scored positively, increases negatively), (ii) tracking geopolitical risk indices focused on oil-producing regions (e.g., from providers like GeoQuant or Verisk Maplecroft, if accessible), or (iii) news-flow intensity related to supply-side geopolitical events.
The correlations below use these conceptual proxies.
A.2 Copper Latent Factor Correlations
@@
A.3 Crude Oil Latent Factor Correlations